Circular No. 6926

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bu IN NGC 3368 B. E. Schaefer, Yale University, reports the following magnitudes ob-

tained with the 3.5-m WIYN telescope and based on ten Landolt comparison stars: May 23.18 UT, B = 12.34 ± 0.04, V = 11.92 ± 0.03; 26.14, 12.44 ± 0.04, 11.94 ± 0.03. Schaefer remarks: “Using these and other reported magnitudes (IAUC 6899, 6902, 6905, 6921), standard-template fits to the light curves of type-Ia supernovae (Schaefer 1996, Ap.J. 460, L19) suggest that peak magnitudes occurred on May 21.0 ± 0.3, with B = 12.36 ± 0.05, V = 11.93 ± 0.04, and a decline rate in B from peak to 15 days later of 0.95 ± 0.05.”

SUPERNOVA 1998cf IN NGC 3504 S. Ryder, Joint Astronomy Centre, reports that he obtained images in

J, H, and K of the nucleus of NGC 3504 on Apr. 23.3 UT with the U.K. Infrared Telescope (+ IRCAM3). These show SN 1998cf (cf. IAUC 6914) at J = 14.9, H = 14.4, and K = 14.2, indicating that it was already fading in the near-infrared.

XTE J2012+381 G. G. Pooley, Mullard Radio Astronomy Observatory, Cambridge, re-

ports: “Ryle Telescope observations at 15 GHz show that the radio source reported by Hjellming et al. IAUC 6924) is variable at 15 GHz. No detection was made on May 28.2 (rms 2 mJy) or 29.2 UT (rms 0.3 mJy). On May 30.3 and June 2.18, the flux was 1.6 ± 0.5 and 3.2 ± 0.6 mJy (means over 2.5 and 5.6 hr), respectively.”

COMET C/1998 J1 (SOHO) W. Liller, Viña del Mar, Chile, provides the following V magnitudes

(estimated ± 0.04) of the inner coma (square 75′′ × 75′′ aperture): May 28.962 UT, 6.87; 29.957, 6.96; 31.967, 5.87 (saturated image). Nearly all of the brightness increase took place in the inner 20′′ of the coma.

This outburst has also been noted by visual observers. Selected visual m_{1} estimates: May 26.37 UT, 5.1 (R. McNaught, Bugaldie, N.S.W., Australia, naked eye); 29.34, 5.0 (D. A. J. Seargent, The Entrance, N.S.W., naked eye); 30.70, 5.3 (T. Cooper, Bredell, Kempton Park, S. Africa, 10×50 binoculars); June 1.35, 3.7 (G. J. Garradd, Loomberah, N.S.W., naked eye); 1.40, 3.4 (M. Mattiazzo, Wallaroo, S. Australia, 7×50 binoculars).

1998 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green